# Incentivizing Reasoning for Advanced Instruction-Following of Large Language Models

### v4

@@ 4.3 Ablation Study and Analysis
Both Tables 1 and 3 corroborate the positive roles of DeepScaleR in developing reasoning. The increment of maths problems is positively associated with the growth of CoT tokens and thereafter the performance improvement (see Fig. 37), implying that the mathematic reasoning is crucial and supplementary to the general-purpose reasoning. However, given the same training steps, the model with full maths did not converge yet, suggesting that more iterations are required towards optimum. Discussions on maths generalization are in Sec. A.7.3.
@@
The immediate imitation of expert reasoning not only encourages fast-thinkers to earn format rewards but also stabilizes training and fills the gaps of rule-centric rewards. Without proper guidance, cold-start LLMs can only repeat their inherent shallow thinking and consistently receive negative feedback (as confirmed in Table 1). In this case, either model collapse or reward hacking is prone to occur and thereafter causes catastrophic distribution drift.
@@ Appendix A Appendix
In this appendix, we first provide the descriptions about the mathematic symbols used in the manuscript. Then, provide the detailed descriptions about the benchmarks used in our experiments and their evaluation metrics. Furthermore, we provide the details about the preparation of datasets used for training, including the publicly available DeepScaleR dataset [28] and the self-evolved complex instruction dataset mentioned in Sec. 3.2. The detailed implementations are presented in Sec. A.5 of the appendix. Finally, we present more fine-grained results for in-depth analyses.
@@ A.1 Symbol Description
$y_{1:t-1}$ The tokens of $y$ starting from the index $1$ to the index $t-1$
@@ A.2.1 Benchmarks
Benchmark Size Constraint Taxonomy Composition Type Verification and Evaluation
@@ A.2.2 Evaluation Metrics
A.3 Self-Evolved Complex Instruction Dataset
The self-evolved complex instruction dataset contains 13K complex instructions and their associated verifications (either with codes for python execution or with scoring questions for LLM-as-a-Judge) and reference expert responses. We choose the WildChat [34] and the Alpaca [86] as our instruction pool to perform self-evolving. The types of rules and constraints are sourced and collected from the taxonomy defined in IFEval, CELLO, ComplexBench, CFBench, FB-Bench, FollowBench, and InfoBench. Such preparation ensures the diversity of the rule and constraint pool. The type of composition follows the definition in the ComplexBench. The statistics of our self-evolved dataset can be found in Table 6. In consideration of the existing benchmarks, models, and the instruction pool, both the languages of Chinese and English are considered.
@@ A.3 Self-Evolved Complex Instruction Dataset
Table 6: Statistics of our self-evolved complex instruction dataset.
@@ A.3.2 Generation of Rules and Constraints with Verifications
[[[[The end of Instruction]]]]
[[[[The start of Sub-Instructions and Scoring Questions]]]]
@@
[[[[The end of Sub-Instructions and Scoring Questions]]]]
@@
[[[[The start of the Modified Sub-Instructions and Scoring Questions]]]]
@@
[[[[The end of the Modified Sub-Instructions and Scoring Questions]]]]
@@
[[[[The end of Sub-Instructions]]]]
[[[[The start of the All-in-One Instruction]]]]
@@
[[[[The start of the Modified Sub-Instructions]]]]
@@
[[[[The start of the Modified All-in-One Instruction]]]]
@@
We provide examples of the generated instructions via Qwen2.5-72B-Instruct in Figs 13 and 14 respectively for verifications by code execution and LLM-as-a-Judge.
@@
A.3.3 Quality Check for Filtering of the Self-Evolved Instructions
@@
To make it clearer, we provide the models used in different stages of self-evolving (see Table 7).
Table 7: The detailed models used in self-evolving of complex instructions.
@@
Self-Instruct with Rules and Constraints Qwen2.5-72B-Instruct [95]
@@ A.4 DeepScaleR-Preview-Dataset
”problem”: ”Let $a_{n}=6^{n}+8^{n}$. Determine the remainder upon dividing $a_{83}$ by $49$.”,
@@
Because $7|(6+8)$, we only consider $6^{82}-6^{81}8+\ldots-8^{81}6+8^{82}\pmod{7}$
@@ A.5.2 Training
For each experiment on 1.5B, 7B, and 8B models, we use the same Qwen2.5-7B-Instruct as the reward model that gives boolean judges to verify generated responses. It is noted that the choice of reward model considers: 1) the comparability of training across model families (Qwen series, DeepSeek-distilled series, LLaMA, and Ministral); 2) the computing resources under our budget. We believe that the larger, stronger reward model (e.g., Qwen2.5-72B-Instruct, DeepSeek V3) would provide more accurate judgement. To evaluate the competence of the Qwen2.5-7B-Instruct as a LLM judge, we perform both automatic and manual analysis on its scorings. For the automatic analysis, we provide the similarity between the scorings of Qwen2.5-7B-Instruct, Qwen2.5-72B-Instruct, QwQ-32B and those of DeepSeek R1 on 1K randomly sampled generations (see Table 8). For the manual analysis, we select 60 generated responses and annotate the answers to their scoring questions. The accuracy of DeepSeek R1, QwQ-32B, Qwen2.5-7B-Instruct, and Qwen2.5-72B-Instruct is reported in Table 9. It can be observed from Table 8 that compared with the DeepSeek R1, the Qwen2.5-7B-Instruct model indeed achieves a high recall that can pinpoints correct responses. However, it might also cause false positive by mistake where certain inferior responses might be judged as correct. From Table 9, we can see that the average accuracy of Qwen2.5-7B-Instruct is around 68.8%, which is slightly lower than Qwen2.5-72B-Instruct. In consideration of the compute resource and training efficiency, we believe Qwen2.5-7B-Instruct is indeed an acceptable ”proxy” reward model. Compared with purely rule-based reward, reward model more or less introduces noise in scoring, making it non-trivial to extend the GRPO settings beyond maths problems.
@@
For training our baselines with the same 13K self-evolved complex instructions, we conduct experiments with LLaMA Factory [100] and train all models for 10ep. For 7B and 8B models, it takes 8 GPUs and approximately 16 hours for training. For 1.5B models, it takes 4 GPUs and approximately 12 hours for training.
@@ A.6 Empirical Insights from Preliminary Studies
LLaMA3.3-70B-Instruct I/O 91.50 80.20 72.67 82.12 62.20 88.82 88.49 80.85
LLaMA3.3-70B-Instruct CoT 72.46 72.00 62.67 70.79 39.27 81.29 83.38 68.83(-12.01%)
LLaMA3.3-70B-Instruct SDC 72.64 76.70 66.67 77.42 49.34 85.18 88.49 73.77(-7.07%)
@@ A.7.1 Detailed Results on Complex Instruction Benchmarks
In this section, we provide the detailed results of Table 1 in each benchmark: IFEval (Table 15), CELLO (Table 16), CFBench (Table 17), ComplexBench (Table 18), FBBench (Table 19), FollowBench (Tables 20 and 21), and InfoBench (Table 22). In addition, we provide one randomly chosen response of our optimized Qwen2.5-7B for each benchmark: IFEval (Fig. 25), CELLO (Fig. 26), CFBench (Fig. 27), ComplexBench (Fig. 28), FBBench (Fig. 29), FollowBench (Figs. 30 and 31), and InfoBench (Fig. 32).
@@
The CELLO dataset, to a certain degree, is not discriminative enough for revealing the strength of reasoning as the DeepSeekR1 and QwQ perform quite similar to Qwen2.5-7B-Instruct, LLaMA3.1-8B-Instruct, and Ministral-8B-Instruct. However, all the 1.5B models achieved gains especially on complex instructions.
@@
We find that the ISR results get improved for all models except the degraded LLaMA, suggesting that the satisfaction of atomic sub-instructions indeed gets improved.
@@
It is noted that except DeepSeek-Qwen7B and the degraded LLaMA, the performance of all models on Selection and Selection and Chain is improved, implying that the deep reasoning truly boosts existing LLMs on instructions with sophisticated composition types.
@@ A.7.2 Generalization on Multi-Purpose Benchmarks
  • The LLaMA model, due to its model collapse during the RL process, fail to deliver effective answers that can be successfully parsed in post-processing, suggesting that it encounters severe degeneration.
@@ A.7.3 Generalization on Maths Benchmarks
  • With respect to the model family, we find that Qwen, LLaMA, and Ministral all develop mathematic reasoning that leads to improved results.
  • Especially for LLaMA3.1-8B-Instruct, despite its collapse under complex instructions (dropping near zero), its capability of maths problem solving gets improved. Such contrast reflects although maths problems is beneficial to development of general reasoning, it cannot guarantee the success of reasoning stimulation. We believe the pre-trained multi-lingual (e.g., Chinese) and multi-task (e.g., role-playing) knowledge of base models is also indispensable.
@@ A.8 Limitations and Future Work
LLaMA3.1-8B-Instruct I/O 79.09 73.01 81.18 77.63 84.53
LLaMA3.1-8B-Instruct CoT 62.94 57.49 65.83 60.44 67.99
LLaMA3.1-8B-Instruct SDC 80.60 74.31 81.77 80.22 86.09
LLaMA3.1-8B-Instruct SFT 79.30 74.12 81.77 77.26 84.05
LLaMA3.1-8B-Instruct Ours 19.10 12.56 24.82 13.49 25.53
@@
LLaMA3.1-8B-Instruct I/O 75.33 75.40 73.10 77.70 61.40 69.20 76.70 79.90 83.30 90.20 61.00 81.10 73.20 77.20
LLaMA3.1-8B-Instruct CoT 65.44 66.10 58.20 69.70 66.10 42.20 76.00 79.50 86.00 86.60 30.40 58.80 67.30 63.80
LLaMA3.1-8B-Instruct SDC 76.02 73.30 73.10 73.50 66.00 63.10 78.00 84.60 86.50 90.50 68.70 79.50 73.40 78.30
LLaMA3.1-8B-Instruct SFT 71.10 67.00 56.90 69.40 62.10 56.40 78.50 81.70 87.50 87.10 57.60 78.10 70.30 71.80
LLaMA3.1-8B-Instruct Ours 4.50 0.30 0.70 0.00 4.39 3.40 3.30 2.90 20.40 13.80 0.00 0.00 6.30 2.90
@@
LLaMA3.1-8B-Instruct I/O 79.00 69.00 55.00 29.00 63.00 46.00 74.00 42.00 55.00 57.00
LLaMA3.1-8B-Instruct CoT 70.00 62.00 44.00 23.00 51.00 37.00 66.00 33.00 44.00 47.66
LLaMA3.1-8B-Instruct SDC 81.00 72.00 56.00 32.00 63.00 47.00 76.00 44.00 55.00 58.33
LLaMA3.1-8B-Instruct SFT 80.00 68.00 51.00 25.00 59.00 40.00 74.00 38.00 50.00 54.00
LLaMA3.1-8B-Instruct Ours 5.00 4.00 0.00 0.00 0.00 0.00 4.00 0.00 0.00 1.33
@@
LLaMA3.1-8B-Instruct I/O 81.89 68.05 70.57 69.96 56.22 56.69 52.08 55.81 64.89 52.35 58.62 69.11
LLaMA3.1-8B-Instruct CoT 72.89 40.83 40.28 40.41 64.81 55.58 49.06 51.82 47.33 35.88 41.61 56.54
LLaMA3.1-8B-Instruct SDC 83.33 64.20 61.89 62.45 63.52 59.63 59.62 58.72 53.44 54.12 53.78 68.74
LLaMA3.1-8B-Instruct SFT 80.81 70.12 65.57 66.67 57.94 50.61 44.91 48.75 39.69 43.53 41.61 65.24
LLaMA3.1-8B-Instruct Ours 5.40 1.18 1.60 1.50 1.28 1.31 0.37 0.92 0.00 0.29 0.14 2.71
@@
LLaMA3.1-8B-Instruct I/O 54.21 39.63 46.92
LLaMA3.1-8B-Instruct CoT 45.15 19.53 32.34
LLaMA3.1-8B-Instruct SDC 45.62 31.10 38.36
LLaMA3.1-8B-Instruct SFT 46.11 34.21 40.16
LLaMA3.1-8B-Instruct Ours 12.91 1.37 7.14
@@
LLaMA3.1-8B-Instruct I/O 3.20 83.04 72.00 70.97 68.01 61.92 83.04 75.22 76.56 74.62 73.29
LLaMA3.1-8B-Instruct CoT 2.32 69.88 66.95 60.15 52.42 52.13 69.88 71.71 67.71 64.93 65.57
LLaMA3.1-8B-Instruct SDC 3.10 80.42 79.75 70.18 63.81 61.43 80.42 82.58 76.22 72.34 72.31
LLaMA3.1-8B-Instruct SFT 2.86 77.93 69.05 63.20 60.72 51.99 77.93 72.88 71.19 69.68 65.11
LLaMA3.1-8B-Instruct Ours 0.11 5.24 (8.53) (12.27) (12.97) (9.64) 5.24 (7.14) (10.94) (8.52) (7.54)
@@
LLaMA3.1-8B-Instruct I/O 2.42 72.78 69.32 59.91 53.09 45.50 72.78 74.08 68.07 66.47 61.59
LLaMA3.1-8B-Instruct CoT 1.76 63.68 57.71 55.34 40.75 36.48 63.68 65.92 65.51 57.75 52.06
LLaMA3.1-8B-Instruct SDC 2.56 73.07 68.78 63.95 58.60 57.56 73.07 72.67 72.21 69.13 69.24
LLaMA3.1-8B-Instruct SFT 2.12 69.88 56.38 52.81 50.08 46.96 69.88 63.37 66.76 65.22 61.52
LLaMA3.1-8B-Instruct Ours 0.54 1.74 2.45 (0.18) 2.73 2.69 1.74 3.88 2.99 6.90 7.66
@@
LLaMA3.1-8B-Instruct I/O 84.93 85.32 85.20 69.05 43.55 56.40 85.20
LLaMA3.1-8B-Instruct CoT 80.58 76.73 77.91 65.48 36.29 51.00 77.91
LLaMA3.1-8B-Instruct SDC 84.93 82.82 83.47 69.05 42.74 56.00 83.47
LLaMA3.1-8B-Instruct SFT 82.46 81.86 82.04 65.87 35.89 51.00 82.04
LLaMA3.1-8B-Instruct Ours 2.75 2.63 2.67 0.40 0.40 0.40 2.67
@@
(a) Response length.
(b) Total length.
(c) SFT loss.
(d) KL penalty.
(e) Format reward.
(f) Accuracy reward.
(g) Policy loss.
(h) Return.
Figure 35: Training dynamics across model families: Qwen2.5-7B, LLaMA3.1-8B, and Ministral-8B (best viewed magnified).

### v5

@@ 4.3 Ablation Study and Analysis
Both Tables 1 and 4 corroborate the positive roles of DeepScaleR in developing reasoning. The increment of maths problems is positively associated with the growth of CoT tokens and thereafter the performance improvement (see Fig. 36), implying that the mathematic reasoning is crucial and supplementary to the general-purpose reasoning. However, given the same training steps, the model with full maths did not converge yet, suggesting that more iterations are required towards optimum. Discussions on maths generalization are in Sec. A.7.3.
@@
The immediate imitation of expert reasoning not only encourages fast-thinkers to earn format rewards but also stabilizes training and fills the gaps of rule-centric rewards. Without proper guidance, The instructed models (i.e., non-reasoning model) are not pretrained on reasoning corpus in general domains and therefore they exhibit shallow reasoning. In this case, we perform behavior cloning on instructed models so that 1) the adherence to the expected format allows successful parsing and grading of answers at an early stage; 2) the deep reasoning on complex instructions distilled from DeepSeek R1 can be imitated; 3) the potential reward hacking can be mitigated to prevent over-optimization.
Table 5: Generalization on OOD constraints of IFBench.
Model Method
prompt-level
strict
instruction-level
strict
prompt-level
loose
instruction-level
loose
Avg.
Qwen2.5-1.5B-Instruct I/O 15.64 17.01 18.36 20.29 17.82
Qwen2.5-1.5B-Instruct Ours 17.68 19.4 20.06 22.68 19.95(+2.13%)
DeepSeek-Qwen-1.5B I/O 8.80 11.64 12.92 15.82 12.29
DeepSeek-Qwen-1.5B Ours 12.92 14.02 15.64 16.71 14.82(+2.53%)
DeepScaleR-1.5B I/O 11.22 12.23 15.3 17.91 14.16
DeepScaleR-1.5B Ours 12.58 14.02 17.00 18.80 15.60(+1.44%)
Qwen2.5-7B-Instruct I/O 28.23 29.85 31.63 33.43 30.78
Qwen2.5-7B-Instruct Ours 29.82 30.76 32.27 35.43 32.07(+1.29%)
Ministral-8B-Instruct I/O 16.66 17.31 23.12 24.47 20.39
Ministral-8B-Instruct Ours 20.74 23.88 28.23 31.94 26.19(+5.77%)
DeepSeek-Qwen7B I/O 13.6 14.62 19.72 22.08 17.50
DeepSeek-Qwen7B Ours 20.06 22.38 25.17 27.46 23.77(+6.27%)
4.4 Generalization Study on OOD Constraints
To test if the proposed RAIF can generally improve the performance of LLMs on unseen constraints, we follow Tülu 3 Lambert et al. (2024) to conduct experiments on brand-new complex instructions that differ from existing benchmarks. Specifically, we use the most recently proposed IFBench Pyatkin et al. (2025) as OOD evaluation benchmark because it contains 58 new, diverse, and challenging constraints. There exists no possibility of data contamination with respect to the training and testing data of the present study.
As shown in Table 5, our RAIF achieves consistent improvement on the instruction following capabilities across model sizes and families, demonstrating the generalizability on new constraints. Tülu 3 Lambert et al. (2024) reports that their optimized LLMs overfit IFEval and exhibit performance degradation on IFBench after RLVR. This contradicts with the finding that our RAIF does NOT overfit the constraints of IFEval, which highlights our advantages over existing RLVR methods. We believe such generalization comes from: 1) diversity of the evolved complex instructions with various constraints and composition structures, and 2) the universal application of deep reasoning that benefit analysis and decomposition of instructions.
@@ Appendix A Appendix
In this appendix, we first provide the descriptions about the mathematic symbols used in the manuscript. Then, provide the detailed descriptions about the benchmarks used in our experiments and their evaluation metrics. Furthermore, we provide the details about the preparation of datasets used for training, including the publicly available DeepScaleR dataset [29] and the LLM-evolved complex instruction dataset mentioned in Sec. 3.2. The detailed implementations are presented in Sec. A.5 of the appendix. Finally, we present more fine-grained results for in-depth analyses.
@@ A.1 Symbol Description
$y_{1:t-1}$ The tokens of $y$ starting from the index 111 to the index $t-1$
@@ A.2.1 Benchmarks
Benchmark Size
Constraint
Taxonomy
Composition Type Verification and Evaluation
@@ A.2.2 Evaluation Metrics
A.3 LLM-Evolved Complex Instruction Dataset
The LLM-evolved complex instruction dataset contains 13K complex instructions and their associated verifications (either with codes for python execution or with scoring questions for LLM-as-a-Judge) and reference expert responses. We choose the WildChat [35] and the Alpaca [88] as our instruction pool to perform LLM-based evolving. The types of rules and constraints are sourced and collected from the taxonomy defined in IFEval, CELLO, ComplexBench, CFBench, FB-Bench, FollowBench, and InfoBench. Such preparation ensures the diversity of the rule and constraint pool. The type of composition follows the definition in the ComplexBench. The statistics of our LLM-evolved dataset can be found in Table 8. In consideration of the existing benchmarks, models, and the instruction pool, both the languages of Chinese and English are considered.
@@ A.3 LLM-Evolved Complex Instruction Dataset
Table 8: Statistics of our LLM-evolved complex instruction dataset.
@@ A.3.2 Generation of Rules and Constraints with Verifications
[[[The end of Instruction]]]
[[[The start of Sub-Instructions and Scoring Questions]]]
@@
[[[The end of Sub-Instructions and Scoring Questions]]]
@@
[[[The start of the Modified Sub-Instructions and Scoring Questions]]]
@@
[[[The end of the Modified Sub-Instructions and Scoring Questions]]]
@@
[[[The end of Sub-Instructions]]]
[[[The start of the All-in-One Instruction]]]
@@
[[[The start of the Modified Sub-Instructions]]]
@@
[[[The start of the Modified All-in-One Instruction]]]
@@
We provide examples of the generated instructions in Figs 13 and 14 respectively for verifications by code execution and LLM-as-a-Judge.
@@
A.3.3 Quality Check for Filtering of the LLM-Evolved Instructions
@@
To make it clearer, we provide the models used in different stages of LLM-based evolving (see Table 9).
Table 9: The detailed models used in LLM-based evolving of complex instructions.
@@
Self-Instruct with Rules and Constraints DeepSeek V3 [104]
@@ A.4 DeepScaleR-Preview-Dataset
”problem”: ”Let $a_{n}=6^{n}+8^{n}$. Determine the remainder upon dividing $a_{83}$ by 494949.”,
@@
Because 7|(6+8)7|(6+8)7 | ( 6 + 8 ), we only consider $6^{82}-6^{81}8+\ldots-8^{81}6+8^{82}\pmod{7}$
@@ A.5.2 Training
For each experiment on 1.5B, 7B, and 8B models, we use the same Qwen2.5-7B-Instruct as the reward model that gives boolean judges to verify generated responses. It is noted that the choice of reward model considers: 1) the comparability of training across model families (Qwen series, DeepSeek-distilled series, and Ministral); 2) the computing resources under our budget. We believe that the larger, stronger reward model (e.g., Qwen2.5-72B-Instruct, DeepSeek V3) would provide more accurate judgement. To evaluate the competence of the Qwen2.5-7B-Instruct as a LLM judge, we perform both automatic and manual analysis on its scorings. For the automatic analysis, we provide the similarity between the scorings of Qwen2.5-7B-Instruct, Qwen2.5-72B-Instruct, QwQ-32B and those of DeepSeek R1 on 1K randomly sampled generations (see Table 10). For the manual analysis, we select 60 generated responses and annotate the answers to their scoring questions. The accuracy of DeepSeek R1, QwQ-32B, Qwen2.5-7B-Instruct, and Qwen2.5-72B-Instruct is reported in Table 11. It can be observed from Table 10 that compared with the DeepSeek R1, the Qwen2.5-7B-Instruct model indeed achieves a high recall that can pinpoints correct responses. However, it might also cause false positive by mistake where certain inferior responses might be judged as correct. From Table 11, we can see that the average accuracy of Qwen2.5-7B-Instruct is around 68.8%, which is slightly lower than Qwen2.5-72B-Instruct. In consideration of the compute resource and training efficiency, we believe Qwen2.5-7B-Instruct is indeed an acceptable ”proxy” reward model. Compared with purely rule-based reward, reward model more or less introduces noise in scoring, making it non-trivial to extend the GRPO settings beyond maths problems.
@@
For training our baselines with the same 13K LLM-evolved complex instructions, we conduct experiments with LLaMA Factory [105] and train all models for 10ep. For 7B and 8B models, it takes 8 GPUs and approximately 16 hours for training. For 1.5B models, it takes 4 GPUs and approximately 12 hours for training.
@@ A.7.1 Detailed Results on Complex Instruction Benchmarks
In this section, we provide the detailed results of Table 1 in each benchmark: IFEval (Table 17), CELLO (Table 18), CFBench (Table 19), ComplexBench (Table 20), FBBench (Table 21), FollowBench (Tables 22 and 23), and InfoBench (Table 24). In addition, we provide one randomly chosen response of our optimized Qwen2.5-7B for each benchmark: IFEval (Fig. 25), CELLO (Fig. 26), CFBench (Fig. 27), ComplexBench (Fig. 28), FBBench (Fig. 29), FollowBench (Figs. 30 and 31), and InfoBench (Fig. 32).
@@
The CELLO dataset, to a certain degree, is not discriminative enough for revealing the strength of reasoning as the DeepSeekR1 and QwQ perform quite similar to Qwen2.5-7B-Instruct, and Ministral-8B-Instruct. However, all the 1.5B models achieved gains especially on complex instructions.
@@
We find that the ISR results get improved for all models suggesting that the satisfaction of atomic sub-instructions indeed gets improved.
@@
It is noted that except DeepSeek-Qwen7B, the performance of all models on Selection and Selection and Chain is improved, implying that the deep reasoning truly boosts existing LLMs on instructions with sophisticated composition types.
@@ A.7.3 Generalization on Maths Benchmarks
  • With respect to the model family, we find that Qwen, and Ministral all develop mathematic reasoning that leads to improved results.